# Benchmark on Drug Target Interaction Modeling from a Structure Perspective

### v1

@@ Abstract
The prediction modeling of drug-target interactions is crucial to drug discovery and design, which has seen rapid advancements owing to deep learning technologies. Recently developed methods, such as those based on graph neural networks (GNNs) and Transformers, demonstrate exceptional performance across various datasets by effectively extracting structural information. However, the benchmarking of these novel methods often varies significantly in terms of hyperparameter settings and datasets, which limits algorithmic progress. In view of these, we conduct a comprehensive survey and benchmark for drug-target interaction modeling from a structure perspective, via integrating tens of explicit (i.e., GNN-based) and implicit (i.e., Transformer-based) structure learning algorithms. To this end, we first unify the hyperparameter setting within each class of structure learning methods. Moreover, we conduct a macroscopical comparison between these two classes of encoding strategies as well as the different featurization techniques that inform molecules’ chemical and physical properties. We then carry out the microscopical comparison between all the integrated models across the six datasets, via comprehensively benchmarking their effectiveness and efficiency. Remarkably, the summarized insights from the benchmark studies lead to the design of model combos. We demonstrate that our combos can achieve new state-of-the-art performance on various datasets associated with cost-effective memory and computation. Our code is available at https://github.com/justinwjl/GTB-DTI/tree/main.
1 Introduction
@@
The prediction modeling of drug-target interactions (DTI) has emerged as an irreplaceable task for efficacious therapeutic interventions. The binding affinity between a drug molecule and its target protein plays a significant role in the design and repurpose of drugs, where a high affinity typically indicates the desired therapeutics, target specificity, long residence, and drug resistance delay [1, 2, 3]. The precise modeling of DTI can expedite the drug discovery process and circumvent the associated cost [4, 5]. Deep learning based frameworks have recently revolutionized this field, enabling more accurate and efficient predictions compared with laboratory experimental methods [6, 7, 8].
@@
Within the deep learning frameworks [9, 10], drugs are commonly represented using the Simplified Molecular Input Line Entry System (SMILES) [11], and proteins are represented as sequences of amino acids. These representations are processed by separate convolutional neural networks (CNNs) [12, 13] and subsequently integrated and processed using a multi-layer perceptron (MLP) for DTI prediction. It is notorious that the reliance on sequence-based representations can result in the loss of structural information, which can potentially compromise the DTI predictive capability. From the drug perspective, molecular structure modeling helps identify the specific binding sites [14], contribute to predicting pharmacokinetic properties [15], and allow conformational flexibility [16].
@@
To address this problem, a number of drug algorithms have been proposed to promote DTI prediction, which can be categorized into explicit and implicit structure learning. First, graph neural networks (GNNs) [17, 18] have been widely adopted to learn the molecular structures, owing to their ability to directly operate on graph-based representations of molecules. By explicitly propagating information through the graph, GNNs can learn node and edge features and thereby capture the structural and functional relationships between atoms and bonds. Second, Transformers, originally focused on natural language processing [19], have also shown promise in biomedical applications [20, 21]. They rely on self-attention mechanisms to implicitly weight the correlations between different parts of the input SMILES, allowing them to capture long-range dependencies and contextual information.
@@
While these techniques contribute to the learning of drug structures, there is still a key knob under-explored: we lack a systematic study to benchmark their effectiveness and efficiency. Without such a standardized benchmark, it is unachievable to offer fair comparisons and subsequently summarize designing philosophy necessary to inform DTI. There have been several surveys and benchmarks on computational methods for DTI prediction [9, 8, 22, 23], which leave out the recent developments of structure learning algorithms and unavoidably fail to focus on drug structure benchmarking. Moreover, although massive efforts [24, 25, 18] have been made to explore the effectiveness of modeling structural information, they predominantly use their proprietary training hyperparameters, datasets, and evaluation metrics. Due to the various settings, one cannot reach convincing answers whether a configuration of structure encoders and/or featurization methods generally performs well. The complex of DTI classification and regression tasks and datasets complicates the benchmark comparison.
@@
In this study, we introduce GTB-DTI, a comprehensive benchmark customized for GNN and Transformer-based methodologies for DTI prediction. i) We thoroughly examine the implementation details for each category of drug structure learning methods and integrate three widely-used datasets for classification and regression tasks, respectively. Then, we harmonize the sensitive hyperparameters across different methods using a greedy search to identify an optimal sweet spot configuration. The unified setting lays the foundation for a fair and reproducible benchmark. ii) To gain macroscopical insights into the structure encoders and featurization methods, we fix the drug encoder to be either GNN or Transformer-based approaches and benchmark these two strategies in the various settings. We also integrate tens of drug features given their importance to inform molecules’ chemistry and physical properties and evaluate them on the representative datasets. iii) To gain macroscopical insights into nuance between 31 concerned models, we conduct the benchmark studies of their effectiveness on the six datasets with the unified setting. Moreover, we assess the efficiency of each method by measuring peak GPU memory usage, running time, and convergences. iv) The comprehensive study finally provides a number of surprising observations: ❶ The CNN-encoder accompanied with integer features has the close protein embedding performance compared to the Transformer or larger language models, but they are more efficient. ❷ The explicit and implicit structure encoders for drugs exhibit unequal performances across the different datasets, which suggests their hybrid usage for generalization purpose. ❸ Inspired from these insights, we conclude with a model combos that leads us to attaining state-of-the-art (SOTA) regression results and performing similarly to SOTA in the DTI classifications. Our combos further deliver cost-effective memory usage and running time as well as faster convergence, which can serve as new baseline for the following explorations.
@@ 2 Formulations for Drug-target Interaction Modeling
In this research, we focus on the formulations of recently-emerging structure modeling approaches for drug molecules, which could be categorized into explicit methods based on graph neural networks and implicit methods based on transformer. The target proteins are learned by the sophisticated tools of convolutional/recurrent neural networks (CNNs/RNNs) or transformers, after which both the molecules’ and proteins’ embeddings are integrated to facilitate interaction prediction. We will also summarize and benchmark the various widely-adopted molecule features.
@@ 2.1 Graph Neural Networks based Methods
A drug molecule is typically represented as a graph , where and denotes the sets of atoms and chemical bonds, respectively. The classical GNN frameworks involve key processes of aggregating and updating node features, collectively referred to as message passing, which can be mathematically represented as [26, 27]:
@@
where is the feature representation of node at layer , is the feature representation of edge between nodes and , refers to the set of neighboring nodes next to node . Functions and aim to aggregate the neighborhood representations and integrate them together with the nodes features, respectively. Additionally, and are feature mapping functions, parameterized by and , respectively. The molecule’s representation can be derived using READOUT function, which processes on the set of vertex features at the last layer.
@@
Graph Convolutional Networks (GCN). Given a molecule with atoms, the adjacency matrix indicates its connectivity, with if atom is adjacent to atom , and otherwise. Considering the self-connection of atoms, we have . Let denote the initial atom feature matrix. GCN [28] models the message passing as follows:
@@
where is the node feature matrix at layer , starting with . Matrix represents the learnable weights for layer , denotes a non-linear activation function, e.g., ReLU, and is a diagonal degree matrix of . A couple of pioneering works have leveraged GCN to facilitate drug-protein interaction prediction [29, 30, 31, 32]. For example, DeepGLSTM [29] uses mixture-of-depths GCNs to capture drug representations from different scales. CPI [31] considers cross-atom distance and introduces the concept of r-radius subgraphs [33], using r-radius vertices and edges to redefine the structure of graphs.
@@
Graph Isomorphism Networks (GIN). GIN excels in learning distinct graph features by approximating the Weisfeiler-Lehman test, enabling it to distinguish a wide range of graph structures [34]. The message passing process at the -th layer is of the following form:
@@
where is a multi-layer perceptron that parameterizes the update function, and is a learnable parameter. We benchmark several GIN-based drug-target interaction modeling methods. GraphCPI [35] and GraphDTA [18] adopt GIN-based models with batch normalization to obtain the drug representation. SubMDTA [36] uses subgraph’s generation task and contrastive learning to pretrain a molecular graph encoder with multiple GIN layers for further prediction.
@@
Graph Attention Networks (GAT). Unlike fixed-weight aggregation, GAT [37] employs an attention mechanism to determine neighborhood importance and learn the node embeddings as:
@@ 2.2 Transformer-based Methods
Besides the graph representation, drugs could also be decorated as SMILES strings [39] and encoded similarly to natural language processing. Specifically, after tokenizing SMILES strings, Transformer model utilizes multi-head attention to model the interactions between different segments of the input and obtain the molecular representations. Positional encodings are also integrated to preserve the sequence order, enhancing the model’s ability to process sequential information effectively. We review and benchmark two typical types of attention mechanisms used for molecular representations.
Self-Attention [20, 40, 41]. Self-attention computes a weighted sum of all input values based on their relevance to each other. Considering an embedding of SMILES sequence at a specific transformer layer, where and are token length and dimension, respectively, the attention is calculated by . and are projections of the input matrix . Multi-head attention combines these projections across different subspaces for a more detailed analysis. Following by normalization and feed-forward neural networks, the SMILES embedding is updated as and the output from the last layer is treated as molecular representations. Transformer encoders like MolTrans [20] and FOTFCPI [41] are adopted to enhance sub-structure embeddings in proteins and drugs.
Cross Attention [42, 40]. Cross-attention is designed to capture the interaction between the drug and protein sequences, with the query matrix derived from one sequence and the key and value matrices from another. This mechanism is particularly useful in integrating hybrid representations such as drug graphs and SMILES [38], as well as drugs and proteins [32, 42].
@@ 2.3 Feature Processing Methods
Sequence Processing Methods. Both drugs and proteins are input as strings of ASCII characters, whose features can be extracted using statistical solutions. Integer encoding [18] simply converts the string to a sequence of integers, which assigns an integer to each character. The N-gram [43] captures the statistical dependencies between characters in an input string. Specifically, a 3-gram model breaks down a sequence into , analyzing the relationship between adjacent characters.
Drug-unique Featurization Methods. The additional chemical properties and structural details of SMILES strings are often considered to gain a more comprehensive understanding. Extended-Connectivity Fingerprints (ECFP) [44, 45], involves generating unique identifiers for atoms based on their local chemical environment and iteratively updating these through a hash function to capture a broader molecular context, ultimately producing a set of fingerprints that represent the molecule’s overall structure. Another approach, RDKit, is used to convert SMILES into molecular graphs [46, 18], where nodes represent the physical and chemical properties of molecules, and bonds are represented by an adjacency matrix. For example, atomic properties such as atom type, degree, and hydrogen information (like the number of explicit hydrogens) are all crucial for constructing a graph. More detailed properties can be found in Appendix F.
Embedding Featurization Methods. Embedding methods are used to translate these discrete sequences into continuous embedding spaces. Notably, Smi2Vec [47] and Prot2Vec [48] convert discrete tokens of drug SMILES and protein sequences into vectors that encapsulate semantic and syntactic similarities, effectively grouping similar tokens together in vector space. Additionally, pretrained language models [24, 49] are increasingly utilized to leverage large-scale learned patterns, fine-tuned to analyze complex protein data representations effectively.
@@ 3 A Fair Benchmark Platform Setup
Benchmark Model and Dataset Selection. From the perspective of reproducibility, we restrict our analysis to models for which the source code has been publicly released. To enhance the comprehensiveness, credibility, and sophistication of our benchmark, we conduct experiments on more than 30 models, including both GNN-based and Transformer-based methods. These models are derived from papers spanning the years 2018 to 2024.
We run these models on 6 frequently evaluated datasets including both binary interaction classification and continuous affinity regression. For the classification aspect, we utilize datasets including Human [50], Caenorhabditis elegans (C. elegans) [31], and DrugBank [51]. For regression, we employ the Davis [52], KIBA [53], and BindingDB [54] with dissociation constant (Kd) measures datasets, as processed in [22]. The statistical details of these models and datasets are presented in Section B and Table 4 of Appendix, respectively.
Unifying Hyperparameter Configuration. Given the critical role of hyperparameters in achieving optimal performance, we perform a detailed review of the hyperparameters associated with the selected models in Section E of the Appendix. There is significant variability in the hyperparameters across different models, making it unfair to conduct comparisons directly. To achieve equitable comparisons between varied models, we select two representative approaches from both the GNN-based and Transformer-based categories, i.e., GraphDTA [18], GraphCPI [35], MRBDTA [55], and TransformerCPI [21], to perform a greedy hyperparameter search to find their sweet spot for classification and regression tasks, respectively. For the search space of hyperparameters, we mainly focus on the influence of batch size (BS), learning rate (LR), and dropout rate (DR), as these are the common hyperparameters utilized by all models. Additionally, we standardize the hyperparameters for epochs, weight decay, and the choice of optimizer, setting a consistent 1000 epochs for GNN-based methods and 300 epochs for Transformer-based methods, with a weight decay of 0 and the Adam optimizer for all models according to Table 5. We illustrate the selected results for the metrics MSE and CI for the regression task, along with AUC-ROC and accuracy for the classification task, in Fig. 1 and the results of all metrics in Table 6. In all experiments, we employ the five-fold cross-validation method with a random split to evaluate all different methods and report the averaged results.
We observe that different models exhibit distinct preferences for hyperparameters. Taking into account various models and metrics, we recommend the configuration as the sweet point hyperparameter configuration for the GNN-based model. Similarly, for the Transformer-based model, we suggest . We strictly follow it in the following experiments.
4 A Macroscopical Benchmark on Encoder and Featurization Strategies

### v2

# Benchmark on Drug Target Interaction Modeling from a Drug Structure Perspective

@@ Abstract
The prediction modeling of drug-target interactions is crucial to drug discovery and design, which has seen rapid advancements owing to deep learning technologies. Recently developed methods, such as those based on graph neural networks (GNNs) and Transformers, demonstrate exceptional performance across various datasets by effectively extracting structural information. However, the benchmarking of these novel methods often varies significantly in terms of hyperparameter settings and datasets, which limits algorithmic progress. In view of these, we conducted a comprehensive survey and benchmark for drug-target interaction modeling from a structural perspective via integrating tens of explicit (i.e., GNN-based) and implicit (i.e., Transformer-based) structure learning algorithms. We conducted a macroscopical comparison between these two classes of encoding strategies as well as the different featurization techniques that inform molecules’ chemical and physical properties. We then carry out the microscopical comparison between all the integrated models across the six datasets via comprehensively benchmarking their effectiveness and efficiency. To ensure fairness, we investigate model performance under individually optimized configuration. Remarkably, the summarized insights from the benchmark studies lead to the design of model combos. We demonstrate that our combos can achieve new state-of-the-art performance on various datasets associated with cost-effective memory and computation.
11footnotetext: These authors contributed equally to this work.1 Introduction
@@
The prediction modeling of drug-target interactions (DTI) has emerged as an irreplaceable task for efficacious therapeutic interventions. The binding affinity between a drug molecule and its target protein plays a significant role in the design and repurposing of drugs, where a high affinity typically indicates the desired therapeutics, target specificity, long residence, and drug resistance delay (hughes2011principles; copeland2006drug; swinney2004biochemical). The precise modeling of DTI can expedite the drug discovery process and circumvent the associated cost (Ashburn2004; Strittmatter2014). Deep learning-based frameworks have recently revolutionized this field, enabling more accurate predictions and accelerating the discovery of new compounds by guiding laboratory experiments more efficiently (wen2017deep; abbasi2021deep; huang2020deeppurpose).
@@
Within deep learning frameworks (ozturk2018deepdta; ozturk2019widedta), drugs are commonly represented using the Simplified Molecular Input Line Entry System (SMILES)(Smiles), and proteins are represented as sequences of amino acids. These representations are typically processed using various neural network architectures, such as convolutional neural networks (CNNs) (krizhevsky2017imagenet; he2016deep), recurrent neural networks (RNNs), Transformers, and so on, before being integrated and processed by a multi-layer perceptron (MLP) for DTI prediction. It is notorious that the reliance on sequence-based representations can result in the loss of structural information, which can potentially compromise the DTI predictive capability. From the drug perspective, molecular structure modeling helps identify the specific binding sites (ma2011molecular), contributes to predicting pharmacokinetic properties (ekins2007silico), and allows conformational flexibility (karplus2005molecular).
@@
To address this problem, a number of drug algorithms have been proposed to promote DTI prediction, which can be categorized into explicit and implicit structure learning. First, graph neural networks (GNNs) (kipf2016semi; nguyen2021graphdta) have been widely adopted to learn the molecular structures, owing to their ability to directly operate on graph-based representations of molecules. By explicitly propagating information through the graph, GNNs can learn node and edge features and thereby capture the structural and functional relationships between atoms and bonds. Second, Transformers, originally focused on natural language processing (Transformer), have also shown promise in biomedical applications (huang2021moltrans; chen2020Transformercpi). They rely on self-attention mechanisms to implicitly weight the correlations between different parts of the input SMILES, allowing them to capture long-range dependencies and contextual information.
@@
While these techniques contribute to the learning of drug structures, there is still a key knob under-explored: we lack a systematic study to benchmark their effectiveness and efficiency. Without such a standardized benchmark, it is unachievable to offer fair comparisons and subsequently summarize the design philosophy necessary to inform DTI. There have been several surveys and benchmarks on computational methods for DTI prediction(ozturk2018deepdta; huang2020deeppurpose; huang2021therapeutics; xu2022peer), which leave out the recent developments of structure learning algorithms and unavoidably fail to focus on drug structure benchmarking. Moreover, although massive efforts (bal2024pgraphdta; zhu2023tdgraph; nguyen2021graphdta) have been made to explore the effectiveness of modeling structural information, they predominantly use their proprietary training hyperparameters, datasets, and evaluation metrics. Due to the various settings, one cannot reach convincing answers as to whether a configuration of structure encoders and/or featurization methods generally performs well. The complex of DTI classification and regression tasks and datasets complicates the benchmark comparison.
@@
In this study, we introduce GTB-DTI, a comprehensive benchmark customized for GNN and Transformer-based methodologies for DTI prediction. I) We thoroughly examine the implementation details for each category of drug structure learning methods and integrate three widely used datasets for classification and regression tasks, respectively. We use the optimal hyperparameters reported in their papers to lay the foundation for a fair and reproducible benchmark. II) To gain macroscopical insights into the structure encoders and featurization methods, we fix the drug encoder to be either GNN or Transformer-based approaches and benchmark these two strategies in the various settings. We also integrate tens of drug features given their importance to inform molecules’ chemistry and physical properties and evaluate them on the representative datasets. III) To gain macroscopical insights into nuance between 31 concerned models, we conduct the benchmark studies of their effectiveness on the six datasets. Moreover, we assess the efficiency of each method by measuring peak GPU memory usage, running time, and convergences. IV) The comprehensive study finally provides a number of surprising observations: i) The CNN encoder accompanied by integer features has close protein embedding performance compared to the Transformer or larger language models, but they are more efficient. ii) The explicit and implicit structure encoders for drugs exhibit unequal performances across the different datasets, which suggests their hybrid usage for generalization purposes. iii) Inspired by these insights, we conclude with a model combo that leads us to attaining state-of-the-art (SOTA) regression results and performing similarly to SOTA in the DTI classifications. Our combos further deliver cost-effective memory usage and running time as well as faster convergence, which can serve as a new baseline for the following explorations.
@@ 2 Formulations for Drug-target Interaction Modeling
In this research, we focus on the formulations of recently emerging structure modeling approaches for drug molecules, which could be categorized into explicit methods based on graph neural networks and implicit methods based on Transformers. The target proteins are learned by the sophisticated tools of convolutional/recurrent neural networks (CNNs/RNNs) or Transformers, after which both the molecules’ and proteins’ embeddings are integrated to facilitate interaction prediction. We will also summarize and benchmark the various widely adopted molecule features.
@@ 2.1 Graph Neural Networks based Methods
A drug molecule is typically represented as a graph , where and denote the sets of atoms and chemical bonds, respectively. The classical GNN frameworks involve key processes of aggregating and updating node features, collectively referred to as message passing, which can be mathematically represented as follows(scarselli2008graph; duan2022comprehensive):
@@
where is the feature representation of node at layer , is the feature representation of edge between nodes and , refers to the set of neighboring nodes next to node . Functions , aim to aggregate the neighborhood representations and integrate them together with the node features, respectively. Additionally, and are feature mapping functions, parameterized by and , respectively. The molecule’s representation can be derived using READOUT a function that processes the set of vertex features at the last layer.
@@
Graph Convolutional Networks (GCN). Given a molecule with atoms, the adjacency matrix indicates its connectivity, with if atom is adjacent to atom , and otherwise. Considering the self-connection of atoms, we have . Let’s denote the initial atom feature matrix. GCN (gcn) models the message passing as follows:
@@
where is the node feature matrix at layer , starting with . Matrix represents the learnable weights for layer , denotes a non-linear activation function, e.g., ReLU, and is a diagonal degree matrix of . A couple of pioneering works have leveraged GCN to facilitate drug-protein interaction prediction (mukherjee2022deepglstm; tran2022deepnc; tsubaki2019compound; pan2023csdti). For example, DeepGLSTM (mukherjee2022deepglstm) uses mixture-of-depths GCNs to capture drug representations from different scales. CPI (tsubaki2019compound) considers cross-atom distance and introduces the concept of r-radius subgraphs (r-radius), using r-radius vertices and edges to redefine the structure of graphs.
@@
Graph Isomorphism Networks (GIN). GIN excels in learning distinct graph features by approximating the Weisfeiler-Lehman test, enabling it to distinguish a wide range of graph structures (xu2018powerful). The message-passing process at the -th layer is of the following form:
@@
where is a multi-layer perceptron that parameterizes the update function, and is a learnable parameter. We benchmark several GIN-based drug-target interaction modeling methods. GraphCPI (quan2019graphcpi) and GraphDTA (nguyen2021graphdta) adopt GIN-based models with batch normalization to obtain the drug representation. SubMDTA (pan2023submdta) uses a subgraph’s generation task and contrastive learning to pretrain a molecular graph encoder with multiple GIN layers for further prediction.
@@
Graph Attention Networks (GAT). Unlike fixed-weight aggregation, GAT (gat) employs an attention mechanism to determine neighborhood importance and learn the node embeddings as follows:
denotes attention weights, and is a concatenating operation. GraphDTA (nguyen2021graphdta) and AMMVF (wang2023ammvf) leverage the multi-head GAT layers to optimize the atom messaging. They integrate GAT with other architectural modules, such as GCN, facilitating a more comprehensive representation of drugs.
Graph Transformers. Graph Transformers (graphTransformer1; maziarka2020molecule) have emerged as powerful alternatives to traditional graph neural networks (GNNs) for molecular representation learning. Unlike conventional GNNs, which rely on message-passing mechanisms to propagate local node information, Graph Transformers leverage self-attention mechanisms to capture both local and global dependencies more effectively. By integrating Message Passing Networks into Transformer-style architectures, these models enhance expressiveness, enabling more comprehensive encoding of molecular structures. This hybrid approach allows Graph Transformers to preserve structural information while benefiting from the flexibility of attention-based learning.
@@ 2.2 Transformer-based Methods
Besides the graph representation, drugs could also be decorated as SMILES strings (weininger1988smiles) and encoded similarly to natural language processing. Specifically, after tokenizing SMILES strings, the Transformer model utilizes multi-head attention to model the interactions between different segments of the input and obtain the molecular representations. Positional encodings are also integrated to preserve the sequence order, enhancing the model’s ability to process sequential information effectively. We review and benchmark two typical types of attention mechanisms used for molecular representations.
Self-Attention(huang2021moltrans; qian2023mcl; yin2024fotf). Self-attention computes a weighted sum of all input values based on their relevance to each other. Considering an embedding of a SMILES sequence at a specific Transformer layer, where and are token length and dimension, respectively, the attention is calculated by . and are projections of the input matrix . Multi-head attention combines these projections across different subspaces for a more detailed analysis. Followed by normalization and feed-forward neural networks, the SMILES embedding is updated, and the output from the last layer is treated as molecular representations. Transformer encoders like MolTrans (huang2021moltrans) and FOTFCPI (yin2024fotf) are adopted to enhance substructure embeddings in proteins and drugs.
Cross-Attention(kurata2022ican; qian2023mcl). Cross-attention is designed to capture the interaction between the drug and protein sequences, with the query matrix derived from one sequence and the key and value matrices from another. This mechanism is particularly useful in integrating hybrid representations such as drug graphs and SMILES (wang2023ammvf), as well as drugs and proteins (pan2023csdti; kurata2022ican).
@@ 2.3 Feature Processing Methods
Sequence Processing Methods. Both drugs and proteins are input as strings of ASCII characters, whose features can be extracted using statistical solutions. Integer encoding (nguyen2021graphdta) simply converts the string to a sequence of integers, which assigns an integer to each character. The N-gram (10.1093/bioinformatics/bti801) captures the statistical dependencies between characters in an input string. Specifically, a 3-gram model breaks down a sequence into , analyzing the relationship between adjacent characters.
Drug-Unique Featurization Methods. The additional chemical properties and structural details of SMILES strings are often considered to gain a more comprehensive understanding. Extended-Connectivity Fingerprints (ECFP) (doi:10.1021/c160017a018Morgan; doi:10.1021/ci100050tecfp), involves generating unique identifiers for atoms based on their local chemical environment and iteratively updating these through a hash function to capture a broader molecular context, ultimately producing a set of fingerprints that represent the molecule’s overall structure. Another approach, RDKit, is used to convert SMILES into molecular graphs (landrum2006rdkit; nguyen2021graphdta), where nodes represent the physical and chemical properties of molecules, and bonds are represented by an adjacency matrix. For example, atomic properties such as atom type, degree, and hydrogen information (like the number of explicit hydrogens) are all crucial for constructing a graph. More detailed properties can be found in Appendix G.
Embedding Featurization Methods. Embedding methods are used to translate these discrete sequences into continuous embedding spaces. Notably, Smi2Vec (8621313Smi2Vec) and Prot2Vec (10.1371/journal.pone.0141287Prot2Vec) convert discrete tokens of drug SMILES and protein sequences into vectors that encapsulate semantic and syntactic similarities, effectively grouping similar tokens together in vector space. Additionally, pretrained language models (bal2024pgraphdta; lin2022language) are increasingly utilized to leverage large-scale learned patterns, fine-tuned to analyze complex protein data representations effectively.
@@ 3 A Fair Benchmark Platform Setup
Benchmark Model and Dataset Selection. From the perspective of reproducibility, we restrict our analysis to models for which the source code has been publicly released. To enhance the comprehensiveness, credibility, and sophistication of our benchmark, we conduct experiments on more than 30 models, including both GNN-based and Transformer-based methods. These models are derived from papers spanning the years 2018 to 2024. We run these models on 6 frequently evaluated datasets, including both binary interaction classification and continuous affinity regression. For the classification aspect, we utilize datasets including Human (liu2015improving), Caenorhabditis elegans (C. elegans) (tsubaki2019compound), and DrugBank (wishart2008drugbank). For regression, we employ the Davis (davis2011comprehensive), KIBA (tang2014making), and BindingDB datasets (liu2007bindingdb) with dissociation constant (Kd) measures, as processed in huang2021therapeutics. The statistical details of these models and datasets are presented in Appendix B and Table 3, respectively.
Hyperparameter Configuration. Given the critical role of hyperparameters in achieving optimal performance, we perform a systematic review of the hyperparameters for all selected models in Appendix E. To ensure that comparisons across models are equitable, we consider comparing each model using its optimal hyperparameters, as reported in the corresponding perspective papers.
Data Split. We treat each dataset independently to prevent any information leakage that could arise from training a single model on multiple datasets. For duplicate drug–protein pairs in the regression dataset, only the entry with the maximum affinity score is retained. For duplicates in the classification dataset, all entries are removed if conflicting labels are present; otherwise, a single instance is kept. Following data cleaning, each dataset is split into a training set and a test set. We apply k-fold cross-validation on each training set, where each fold consists of a unique training–validation split, and models trained on different folds are completely independent, thereby eliminating any possibility of cross-fold leakage. The training sets are used to fit the models, the validation sets are used to select the best model during training, and the test sets are for final evaluation.
Other Training Details. We train on the training set and see its performance on the validation set at the end of every epoch, and the model that achieves the best validation performance will be saved. After training, we evaluate the saved model on the test set and save their results. We average the final performance metrics across all folds as the final results. Considering the original training epochs, we use 1000 as maximum epochs limitation. To avoid overfitting, we consider an early stop mechanism in training. Given the complexity in the dataset, we use 50 patience for all datasets. We use MSE as an early stopping evaluation metric for regression and F1 for classification. The detailed results are provided in Appendix F.
4 A Macroscopic Benchmark on Encoder and Featurization Strategies